# Generative Bayesian optimization: generative models as acquisition functions

### v1

@@ 3 A general recipe for generative Bayesian optimization
As seen in Section 2, using generative models for BO typically involves training a regression or classification model as an intermediate step to then train the candidate generator. The use of an intermediate model demands additional computational resources and brings in further sources of approximation errors which may hinder performance. Hence, we propose a framework to train the generative model directly from (noisy) observation values. The main idea is to train the model to approximate a target distribution proportional to BO’s acquisition function and then use the learned generative model as a proposal for the next query locations. There are different approaches to do so, some of which have been previously explored in the literature, for specific acquisition functions, such as the probability of improvement (Brookes et al., 2019; Steinberg et al., 2025) and upper confidence bound (Yun et al., 2025). However, we here focus on a general recipe to turn a generative model into a density following any acquisition function that can be expressed as an expected utility.
@@
Consider a likelihood-free BO setting (Song et al., 2022), where we aim to directly learn an acquisition function $a_{t}:{\mathcal{{X}}}\to\mathbb{{R}}$ at every BO round $t\in\{1,\dots,T\}$ based on available data. If our acquisition function takes the form of an expected utility:
@@
given a threshold $\tau_{t}$ for improvement-based utilities, e.g., $\tau_{t}:=\max_{i<t}y_{i}$ or a quantile of the empirical marginal observations distribution (Tiao et al., 2021). A comprehensive summary of typical utility functions for BO can be found in Wilson et al. (2018). The ones listed above, however, we can write directly as a function of the observations. We also use a soft-plus version of EI (sEI) in our experiments, which remains positive at a low value when $y=\tau$.
@@
As an example, consider the case of PI where $a({\boldsymbol{\mathbf{x}}})=\mathbb{E}[\mathbb{I}[y\geq\tau]]=p(y\geq\tau|{\boldsymbol{\mathbf{x}}})$, which has been previously applied to train generative models for black-box optimization via surrogates (Steinberg et al., 2025). Given a sampler for the conditional distribution $p({\boldsymbol{\mathbf{x}}}|y\geq\tau)$, by Bayes rule, we recover the original PI as:
@@
A batch ${\mathcal{{B}}}_{t+1}:=\{{\boldsymbol{\mathbf{x}}}_{t+1,i}\}_{i=1}^{B}$ is sampled from the learned proposal $q_{t}$. We evaluate the utilities $u_{t+1}(y_{t+1,i})$ with the collected observations $y_{t,i}\sim p(y|{\boldsymbol{\mathbf{x}}}_{t+1,i})$, for $i\in\{1,\dots,B\}$, and repeat the cycle up to a given number of iterations $T\in\mathbb{{N}}$. In the following, we describe approaches to formulate general loss functions for learning acquisition functions and how to ensure that the sequence of batches $\{{\mathcal{{B}}}_{t}\}_{t=1}^{\infty}$ asymptotically
@@ 3.1 Preference-based learning
We aim to apply a similar reparameterization trick to the one in DPO to simplify generative BO methods. Note that, for a general classification loss, such as the one in Equation 5, it is not possible to eliminate the partition function resulting from a DPO-like reparameterization without resorting to approximations, which might change the learned model. Hence, we need a pairwise-contrastive objective.
@@
$\ell^{\mathrm{PL}}_{i}(q,\Delta u_{i}):=-\log\sigma\left(\tau\operatorname{sign}(\Delta u_{i})\left(\log\left(\frac{q({\boldsymbol{\mathbf{x}}}_{i,1})}{p_{0}({\boldsymbol{\mathbf{x}}}_{i,1})}\right)-\log\left(\frac{q({\boldsymbol{\mathbf{x}}}_{i,2})}{p_{0}({\boldsymbol{\mathbf{x}}}_{i,2})}\right)\right)\right),$ (17)
where $\Delta u_{i}:=u_{i,1}-u_{i,2}$, as in the DPO formulation, $\tau>0$ is a (optional) temperature parameter and the prior $p_{0}$ can be given by a reference model, either pre-trained or derived from expert knowledge about feasible solutions to the optimization problem (1). Similar to Rafailov et al. (2023), the learned generative model is seeking to approximate:
$p_{u}^{*}({\boldsymbol{\mathbf{x}}}):=\frac{1}{\zeta_{u}}p_{0}({\boldsymbol{\mathbf{x}}})\exp\left(\frac{1}{\tau}\mathbb{E}[u(y)|{\boldsymbol{\mathbf{x}}}]\right),$ (18)
@@
As shown in Chowdhury et al. (2024), the original DPO loss is not robust to preference noise. As in BO, one usually only observes noisy evaluations of the objective function, utility values directly derived from the observation values will also be noisy and correspondingly the sign of their differences as well. Namely, assume there is a small $p_{\mathrm{flip}}\in(0,1/2)$ probability of the preference directions being flipped w.r.t. the sign of the true expected utility:
@@
Chowdhury et al. (2024) showed that the original DPO preference loss is biased in this noisy case, and proposed a robust version of the DPO loss to address this issue as:
@@ 3.2 Divergence-based learning
which we write in a condensed form to avoid notation clutter with $q_{i}=q_{\lfloor i/B\rfloor}$ and $n$ corresponding to the total number of observations up to a given round. The objective above is unbiased and its global optimum can be shown to converge to $p_{u}^{*}$ by an application of standard results from the adaptive importance sampling literature (Delyon & Portier, 2018). A simpler version of this training objective was derived for CbAS (Brookes et al., 2019) using only the last batch for training, which would allow for convergence as the batch size goes to infinity $B\to\infty$. Furthermore, as we will see in our analysis, convergence to $p_{u}^{*}$ is not sufficient to ensure convergence to the global optima of the objective function $f$.
@@ 3.3 Generalizations
In general, we can extend the above framework to use proper scoring rule $S:{\mathcal{{P}}}({\mathcal{{X}}})\times{\mathcal{{X}}}\to\mathbb{{R}}$ (Gneiting & Raftery, 2007) other than the log loss. We can then learn $q$ approximating $p_{u}^{*}$ by minimizing:
@@
Although we leave the exploration of this formulation for future work, it is readily extensible to other types of generative models which may not have densities available in closed form, such as diffusion and flow matching (Lipman et al., 2024), which still provide flexible probabilistic models.
@@ 4 Theoretical analysis
We make a few mild regularity assumptions about the problem setting and the model. Firstly, for the analysis, we assume that both the models $g_{\theta}$ lie in a reproducing kernel Hilbert space (RKHS) ${\mathcal{{H}}}_{k}$ shared with the true log density $g_{*}$, which is such that $p_{u}^{*}({\boldsymbol{\mathbf{x}}})=\exp g_{*}({\boldsymbol{\mathbf{x}}})$. The domain $\mathcal{{X}}$} is assumed to be a compact metric space with main results specialized for the finite discrete setting, i.e., $\lvert{\mathcal{{X}}}\rvert$$\infty$. The model $q_{\theta}({\boldsymbol{\mathbf{x}}})$ is continuously twice differentiable with respect to the parameters ${\theta}\in\Theta$ with bounded second-order derivatives. The individual losses $\ell:\mathbb{{R}}\times\mathbb{{R}}\to\mathbb{{R}}$ are strictly convex and twice differentiable w.r.t. their first argument. In addition, we assume that, at the target $g_{*}$ the individual loss $\ell(m_{i}(g_{*}),z_{i})$ is conditionally sub-Gaussian (Boucheron et al., 2013; Chowdhury & Gopalan, 2017) w.r.t. the data-generating process, basically meaning that the probability distribution of each loss has zero mean and light tails. We also assume that the regularizer $R_{n}:{\mathcal{{Q}}}\to[0,\infty)$ is strongly convex and twice differentiable. We defer the formal assumptions and their discussion to the appendix.
@@
1 and the latter allows us to establish that the model converges to the target $g_{*}$ associated with the target distribution $p_{u}^{*}$ for a given utility function $u$. However, convergence to the target distribution alone does not ensure optimality of the samples ${\boldsymbol{\mathbf{x}}}\sim q_{t}$. The latter is possible by applying results from reward-weighted regression, which shows that training a proposal to maximize $\mathbb{E}_{y\sim p(y|{\boldsymbol{\mathbf{x}}}),{\boldsymbol{\mathbf{x}}}\sim q_{t-1}}[u(y)\log q({\boldsymbol{\mathbf{x}}})]$ yields a sequence of increasing expected rewards $\mathbb{E}[u(y_{t})]\leq\mathbb{E}[u(y_{t+1})]\leq\dots$ (Štrupl et al., 2022, Thm. 4.1). If the maximizer of the sequence of expected utilities converges to the maximizer of the objective function $f$, then the generative BO proposals will concentrate at the true optimum $\boldsymbol{\mathbf{x}}$$_{*}$. Therefore, for KL-based loss functions, one may drop the proposal densities in the importance sampling weights $1/q_{i-1}({\boldsymbol{\mathbf{x}}}_{i})$ to promote this posterior concentration phenomenon, as corroborated by our experimental findings, which generally did not include importance weights. This same concentration of the learned target distribution should also occur with the preference-based loss functions due to the absence of importance-sampling weights.
@@ 5 Experiments
We evaluate the different variants of generative BO (GenBO) on a number of challenging sequence optimization tasks against popular and strong baselines, including CbAS (Brookes et al., 2019), VSD (Steinberg et al., 2025), and LaMBO-2 (Gruver et al., 2023). Details are in Appendix B.
@@ 5.1 Text optimization
For this experiment, we wish to optimize a short sequence (5 letters) to minimize the edit distance to the sequence ALOHA — which is implemented as a poli black-box (González-Duque et al., 2024). Here ${\mathcal{{X}}}={\mathcal{{V}}}^{M}$ where $\mathcal{{V}}$} is the English alphabet, and $M$ is sequence length. Even though this sequence is relatively short, still $|{\mathcal{{X}}}|=|{\mathcal{{V}}}|^{M}>11.8$ million elements. We increase the difficulty by only allowing $|{\mathcal{{D}}}_{0}|=64$ where the minimum edit distance is 4, $B=8$, and $T=10$. We compare GenBO to the classifier guided VSD (Steinberg et al., 2025) and CbAS (Brookes et al., 2019), and to a simple greedy baseline that applies (3) random mutations to its best candidates per-round (González-Duque et al., 2024). For GenBO, VSD and CbAS we use a simple mean-field (independent) categorical proposal distribution, $q$, and a uniform prior, $p_{0}$. VSD and CbAS use a simple embedding and 1-hidden layer MLP classifier for estimating PI. Architectural details and other experimental specifics are given in subsection B.1. Results are summarized in 1(a) where we can see that the random baseline is not able to make much headway and CbAS under-performs due to its limited use of data (last batch only) in retraining. GenBO with the robust preference loss (rPL) and the soft-plus sEI utility is the best performing variant, and initially outperforms VSD, though VSD eventually closes the gap.
@@
Figure 1: Performance of baseline black box optimizers and GenBO variants on the (1(a)) ALOHA, (1(b)) stability, and (1(c)) solvent accessible surface area optimization problems. Algorithms are sorted in descending order of final average regret, and shaded areas correspond to $\pm 1$ standard deviation.
5.2 Protein design
We now consider three protein sequence design tasks where $|{\mathcal{{V}}}|=20$ and we have varying $M$. We again use VSD, CbAS and random mutation as baselines, and add to them the guided diffusion based LaMBO-2 (Gruver et al., 2023). GenBO, VSD and CbAS all share the same generative backbone, which is the causal transformer used in Steinberg et al. (2025); VSD and CbAS also use the same CNN-classifier guide used in that work. We present additional architectural information, and additional experimental details in subsection B.2. We use the black-box implementations in poli for these tasks, and poli-baselines implementations of the random and LaMBO-2 baselines.
The first task we consider is optimization of the Ehrlich functions introduced by Stanton et al. (2024). These are challenging biologically inspired parametric closed-form that explicitly simulate nonlinear (epistatic) effects of sequence on outcome. The outcomes are $y\in\{-1\}\cup[0,1]$ where 1-1 is reserved for infeasible sequences. We use the same protocol as in Steinberg et al. (2025), where we optimize sequences of length $M=\{15,32,64\}$ all with motif lengths of 4, and $|{\mathcal{{D}}}_{0}|=128$, $T=32$ and $B=128$. The results are summarized in Figure 2. We again see that GenBO variants are able to outperform or match the performance of baselines, with KL-based losses yielding the best performance. In higher dimensions with the longest sequence setting, the benefits of the balanced forward KL loss, with its density minimization effect in areas of lower utility, are more evident.
@@ 5.2 Protein design
For our final set of experiments we present two real protein optimization tasks. These experiments have been adapted from Stanton et al. (2022) where the aims are to maximize the stability and solvent accessible surface (SASA) of the proteins respectively. The black-box is the FoldX molecular simulation software (Schymkowitz et al., 2005), and is wrapped by poli (González-Duque et al., 2024). We chose the mRouge red fluorescent protein ($M=228$) as the base protein for the tasks. Both tasks were given $T=20$ rounds, a batch size of $B=64$, and an initial training set of $|{\mathcal{{D}}}_{0}|=88$ as a subset from Stanton et al. (2022). Results are summarized in 1(b) for stability and 1(c) for SASA. All variants of GenBO find the stability task challenging, along with the LaMBO-2 and random baselines. CbAS and especially VSD are better able to stabilize this protein. However, most variants of GenBO far outperform the baselines on the SASA task, and much more rapidly. We believe this is because this task favors extrapolation away from the prior, and the GenBO variants with no prior performed best.

### v2

@@ 3 A general recipe for generative Bayesian optimization
As seen in Section 2, using generative models for BO typically involves training a regression or classification model as an intermediate step to then train the candidate generator. The use of an intermediate model demands additional computational resources and brings in further sources of approximation errors which may hinder performance. Hence, we propose a framework to train the generative model directly from (noisy) observation values. The main idea is to train the model to approximate a target distribution proportional to BO’s acquisition function and then use the learned generative model as a proposal for the next query locations. There are different approaches to do so, some of which have been previously explored in the literature, for specific acquisition functions, such as the probability of improvement (Brookes2019cbas; Steinberg2025variational) and upper confidence bound (Yun2025dibo). However, we here focus on a general recipe to turn a generative model into a density following any acquisition function that can be expressed as an expected utility.
@@
Consider a likelihood-free BO setting (Song2022lfbo), where we aim to directly learn an acquisition function $a_{t}:{\mathcal{{X}}}\to\mathbb{{R}}$ at every BO round $t\in\{1,\dots,T\}$ based on available data. If our acquisition function takes the form of an expected utility:
@@
given a threshold $\tau_{t}$ for improvement-based utilities, e.g., $\tau_{t}:=\max_{i<t}y_{i}$ or a quantile of the empirical marginal observations distribution (Tiao2021). A comprehensive summary of typical utility functions for BO can be found in Wilson2018. The ones listed above, however, we can write directly as a function of the observations. We also use a soft-plus version of EI (sEI) in our experiments, which remains positive at a low value when $y=\tau$.
@@
As an example, consider the case of PI where $a({\boldsymbol{\mathbf{x}}})=\mathbb{E}[\mathbb{I}[y\geq\tau]]=p(y\geq\tau|{\boldsymbol{\mathbf{x}}})$, which has been previously applied to train generative models for black-box optimization via surrogates (Steinberg2025variational). Given a sampler for the conditional distribution $p({\boldsymbol{\mathbf{x}}}|y\geq\tau)$, by Bayes rule, we recover the original PI as:
@@
A batch ${\mathcal{{B}}}_{t+1}:=\{{\boldsymbol{\mathbf{x}}}_{t+1,i}\}_{i=1}^{B}$ is sampled from the learned proposal $q_{t}$. We evaluate the utilities $u_{t+1}(y_{t+1,i})$ with the collected observations $y_{t,i}\sim p(y|{\boldsymbol{\mathbf{x}}}_{t+1,i})$, for $i\in\{1,\dots,B\}$, and repeat the cycle up to a given number of iterations $T\in\mathbb{{N}}$. This process is summarized in Algorithm 1 in the appendix. In the following, we describe approaches to formulate general loss functions for learning acquisition functions and how to ensure that the sequence of batches $\{{\mathcal{{B}}}_{t}\}_{t=1}^{\infty}$ asymptotically concentrates at the optimum $\boldsymbol{\mathbf{x}}$$^{*}$.
@@ 3.1 Preference-based learning
We aim to apply a similar reparameterization trick to the one in DPO to simplify generative BO methods. Note that, for a general classification loss, such as the one in Equation 3, it is not possible to eliminate the partition function resulting from a DPO-like reparameterization without resorting to approximations, which might change the learned model. Hence, we need a pairwise-contrastive training objective.
@@
$\ell^{\mathrm{PL}}_{i}(q,\Delta u_{i}):=-\log\sigma\left(\beta\operatorname{sign}(\Delta u_{i})\left(\log\left(\frac{q({\boldsymbol{\mathbf{x}}}_{i,1})}{p_{0}({\boldsymbol{\mathbf{x}}}_{i,1})}\right)-\log\left(\frac{q({\boldsymbol{\mathbf{x}}}_{i,2})}{p_{0}({\boldsymbol{\mathbf{x}}}_{i,2})}\right)\right)\right),$ (15)
where $\Delta u_{i}:=u_{i,1}-u_{i,2}$, as in the DPO formulation, $\beta>0$ is a (optional) temperature parameter and the prior $p_{0}$ can be given by a reference model, either pre-trained or derived from expert knowledge about feasible solutions to the optimization problem (1). Similar to Rafailov2023dpo, the learned generative model is seeking to approximate:
$p_{u}^{*}({\boldsymbol{\mathbf{x}}}):=\frac{1}{\zeta_{u}}p_{0}({\boldsymbol{\mathbf{x}}})\exp\left(\frac{1}{\beta}\mathbb{E}[u(y)|{\boldsymbol{\mathbf{x}}}]\right),$ (16)
@@
As shown in Chowdhury2024rdpo, the original DPO loss is not robust to preference noise. As in BO, one usually only observes noisy evaluations of the objective function, utility values directly derived from the observation values will also be noisy and correspondingly the sign of their differences as well. Namely, assume there is a small $p_{\mathrm{flip}}\in(0,1/2)$ probability of the preference directions being flipped w.r.t. the sign of the true expected utility:
@@
Chowdhury2024rdpo showed that the original DPO preference loss is biased in this noisy case, and proposed a robust version of the DPO loss to address this issue as:
@@ 3.2 Divergence-based learning
which we write in a condensed form to avoid notation clutter with $q_{i}=q_{\lfloor i/B\rfloor}$ and $n$ corresponding to the total number of observations up to a given round. The objective above is unbiased and its global optimum can be shown to converge to $p_{u}^{*}$ by an application of standard results from the adaptive importance sampling literature (Delyon2018ais). A simpler version of this training objective was derived for CbAS (Brookes2019cbas) using only the last batch for training, which would allow for convergence as the batch size goes to infinity $B\to\infty$. Furthermore, as we will see in our analysis, convergence to $p_{u}^{*}$ is not sufficient to ensure convergence to the global optima of the objective function $f$.
@@ 3.3 Generalizations
In general, we can extend the above framework to use proper scoring rule $S:{\mathcal{{P}}}({\mathcal{{X}}})\times{\mathcal{{X}}}\to\mathbb{{R}}$ (Gneiting2007strictly) other than the log loss. We can then learn $q$ approximating $p_{u}^{*}$ by minimizing:
@@
Although we leave the exploration of this formulation for future work, it is readily extensible to other types of generative models which may not have densities available in closed form, such as diffusion and flow matching (Lipman2024), which still provide flexible probabilistic models.
@@ 4 Theoretical analysis
We make a few mild regularity assumptions about the problem setting and the model. Firstly, for the analysis, we assume that both the models $g_{\theta}$ lie in a reproducing kernel Hilbert space (RKHS) ${\mathcal{{H}}}_{k}$ shared with the true log density $g_{*}$, which is such that $p_{u}^{*}({\boldsymbol{\mathbf{x}}})=\exp g_{*}({\boldsymbol{\mathbf{x}}})$. The domain $\mathcal{{X}}$} is assumed to be a compact metric space with main results specialized for the finite discrete setting, i.e., $\lvert{\mathcal{{X}}}\rvert$$\infty$. The model $q_{\theta}({\boldsymbol{\mathbf{x}}})$ is continuously twice differentiable with respect to the parameters ${\theta}\in\Theta$ with bounded second-order derivatives. The individual losses $\ell:\mathbb{{R}}\times\mathbb{{R}}\to\mathbb{{R}}$ are strictly convex and twice differentiable w.r.t. their first argument. In addition, we assume that, at the target $g_{*}$ the individual loss $\ell(m_{i}(g_{*}),z_{i})$ is conditionally sub-Gaussian (Boucheron2013; Chowdhury2017) w.r.t. the data-generating process, basically meaning that the probability distribution of each loss has zero mean and light tails. We also assume that the regularizer $R_{n}:{\mathcal{{Q}}}\to[0,\infty)$ is strongly convex and twice differentiable. We defer the formal assumptions and their discussion to the appendix.
@@
1 and the latter allows us to establish that the model converges to the target $g_{*}$ associated with the target distribution $p_{u}^{*}$ for a given utility function $u$. However, convergence to the target distribution alone does not ensure optimality of the samples ${\boldsymbol{\mathbf{x}}}\sim q_{t}$. The latter is possible by applying results from reward-weighted regression, which shows that training a proposal to maximize $\mathbb{E}_{y\sim p(y|{\boldsymbol{\mathbf{x}}}),{\boldsymbol{\mathbf{x}}}\sim q_{t-1}}[u(y)\log q({\boldsymbol{\mathbf{x}}})]$ yields a sequence of increasing expected rewards $\mathbb{E}[u(y_{t})]\leq\mathbb{E}[u(y_{t+1})]\leq\dots$ (Strupl2022rwr, Thm. 4.1). If the maximizer of the sequence of expected utilities converges to the maximizer of the objective function $f$, then the generative BO proposals will concentrate at the true optimum $\boldsymbol{\mathbf{x}}$$_{*}$. Therefore, for KL-based loss functions, one may drop the proposal densities in the importance sampling weights $1/q_{i-1}({\boldsymbol{\mathbf{x}}}_{i})$ to promote this posterior concentration phenomenon, as corroborated by our experimental findings, which generally did not include importance weights. This same concentration of the learned target distribution should also occur with the preference-based loss functions due to the absence of importance-sampling weights.
5 Related work
Using generative models for online-optimization is becoming an increasingly popular method for optimization in discrete, mixed discrete-continuous or high dimensional design spaces where classical BO is limited. The following discusses other works applying generative models to BO settings and contrasts them with the reward-model-free active generation framework we propose.
Latent-space BO.
In latent-space BO (LSBO) methods for high-dimensional problems (Gomez-Bombarelli2018; stanton2022accelerating; Gruver2023), one learns a probabilistic representation of a (usually lower-dimensional) manifold of the data jointly with $f$, and performs BO in that space, projecting query points back to the original space at evaluation time. This technique has led to numerous BO methods for high-dimensional and discrete-space optimization (Gomez-Bombarelli2018; Gruver2023; Gonzalez-Duque2024). Learning this latent-space can, however, cause complications. LSBO can suffer poor sample efficiency if the latent-space is learned from the initial training set and then fixed (tripp2020sample). Or poor performance can arise from reconstruction errors between the latent and observation space (lee2025latent). GenBO and other methods like VSD do not suffer from these issues as all inference is done in the observation space. Despite recent advances in the field (Chu2024inversionbased; lee2025latent; Moss2025cowboys), to our knowledge, LaMBO-2 remains state-of-the-art in LSBO for long sequences, like proteins.
Diffusion for BBO.
There has been recent progress in adapting denoising diffusion models to black-box optimization (BBO) tasks, often by learning a model that can be conditioned on observation values, given a dataset of evaluations (Krishnamoorthy23a). Other approaches involve guiding the diffusion process by a given utility function derived from a regression model (Gruver2023; Yun2025dibo). Note, however, that such methodologies are specific to diffusion, whereas we focus on a general approach that can be applied to arbitrary generative models.
LLMs and BO.
Recent work has begun to integrate large language models (LLMs) into BO pipelines, primarily to inject prior knowledge, improve cold-start performance, or offload certain design decisions to a learned policy. Several studies use LLMs as contextual priors over the design space: for example, guiding initialization or proposal generation by leveraging natural-language domain knowledge (liu2024large), or selecting acquisition functions adaptively via an LLM-driven controller (aglietti2025funbo). Other work treats BO as a test-time search tool that an LLM can call to refine or validate its own proposals during inference (agarwal2025searching). Most relevant to our setting is a recent reward-model-free approach for protein engineering (chen2025generalists), which uses LLM preference modeling, akin to DPO, to steer search without an explicit surrogate. This shares the reward-model-free philosophy of GenBO, but differs fundamentally in relying on a general-purpose LLM, whereas GenBO provides a framework for task-specific generative black-box optimization problems with no language interface or pretrained reward structure.
@@ 6 Experiments
We evaluate several variants of generative BO (GenBO) on a number of challenging sequence optimization tasks against popular and strong baselines, including CbAS (Brookes2019cbas), VSD (Steinberg2025variational), and LaMBO-2 (Gruver2023), besides trivial baselines, random mutations and a genetic algorithm (GA) implemented in poli (Gonzalez-Duque2024). As performance measures, we assess the simple regret, $r_{t}:=f({\boldsymbol{\mathbf{x}}}^{*})-\max_{i\leq n_{t}}f({\boldsymbol{\mathbf{x}}}_{i})$, and the cumulative maximum, $\max_{i\leq n_{t}}f({\boldsymbol{\mathbf{x}}}_{i})$, where $n_{t}:=Bt$ is the number of function evaluations up to round $t$. In legend boxes, algorithms are sorted in descending order of final average regret. Shaded areas correspond to $\pm 1$ standard deviation across five different random seeds. Appendix D presents further details about experiment settings and ablations. Table 4 and 5 summarize final results.
@@ 6.1 Text optimization
As a first experiment, we wish to optimize a short sequence (5 letters) to minimize the edit distance to the sequence ALOHA, which is implemented as a poli black-box (Gonzalez-Duque2024poli). Here ${\mathcal{{X}}}={\mathcal{{V}}}^{M}$ where $\mathcal{{V}}$} is the English alphabet, and $M$ is sequence length. Even though this sequence is relatively short, still $|{\mathcal{{X}}}|=|{\mathcal{{V}}}|^{M}>11.8$ million elements. We increase the difficulty by only allowing $|{\mathcal{{D}}}_{0}|=64$ where the minimum edit distance is 4, $B=8$, and $T=10$. We compare GenBO to the classifier guided VSD (Steinberg2025variational) and CbAS (Brookes2019cbas), and to a simple greedy baseline that applies (3) random mutations to its best candidates per-round (Gonzalez-Duque2024poli). For GenBO, VSD and CbAS we use a simple mean-field (independent) categorical proposal distribution, $q$, and a uniform prior, $p_{0}$. VSD and CbAS use a simple embedding and 1-hidden layer MLP classifier for estimating PI. We also varied the threshold $\tau$ annealing schedule. Architectural details and other experimental specifics are given in subsection D.1.
Results are summarized in 1(a). We can see that the random baseline is not able to make much headway and CbAS under-performs due to its limited use of data (last batch only) in retraining. For this experiment, GenBO with the robust preference loss (rPL) and EI-based utilities showed the quickest improvements, whereas PI is able to reach the exact optimum at the end, with VSD eventually also achieving good performance. In Figure 5 (appendix), we present an ablation study on the threshold $\tau_{t}$ annealing scheme we used to balance the exploration-exploitation trade-off for GenBO and PI-based baselines (VSD and CbAS). The plots reveal that this problem generally favors a more exploitative approach by concentrating on higher quantiles of the observations marginal distribution. GenBO was, however, relatively less sensitive to the choice of annealing scheme, as long as the final percentile was set anywhere above 90%, whereas VSD required a generally sharper rise to above the 95% quantile towards the end of the optimization process, favoring original settings suggested by Steinberg2025variational. We also find that in this problem the use of a pre-trained informative prior $p_{0}$ may not bring significant performance advantage, as GenBO variants with no prior (i.e., $p_{0}\propto 1$) performed best. Lastly, we also highlight significant improvements in run time for GenBO, making it on average 3 times faster than VSD (see Table 6 in the appendix) for not needing to fit an intermediate surrogate model.
@@
Figure 1: Performance of baseline black box optimizers and GenBO variants on the ALOHA, stability, and solvent accessible surface area (SASA) optimization problems.
@@
6.2 Protein design
We now consider three protein sequence design tasks where $|{\mathcal{{V}}}|=20$ and we have varying $M$. We again use VSD, CbAS and random mutation as baselines, and add to them the guided diffusion based LaMBO-2 (Gruver2023). GenBO, VSD and CbAS all share the same generative backbone, which is the causal transformer used in Steinberg2025variational; VSD and CbAS also use the same CNN-classifier guide used in that work. We present additional architectural information, and additional experimental details in subsection D.2. We use the black-box implementations in poli for these tasks, and poli-baselines implementations of the random and LaMBO-2 baselines.
The first task we consider is optimization of the Ehrlich functions introduced by stanton2024closed. These are challenging biologically inspired parametric closed-form functions that explicitly simulate nonlinear (epistatic) effects of sequence on outcome. The outcomes are $y\in\{-1\}\cup[0,1]$ where 1-1 is reserved for infeasible sequences. We use the same protocol as in Steinberg2025variational, where we optimize sequences of length $M=\{15,32,64\}$ all with motif lengths of 4, and $|{\mathcal{{D}}}_{0}|=128$, $T=32$ and $B=128$. The results are summarized in Figure 2. We again see that GenBO variants are able to outperform or match the performance of baselines, with KL-based losses yielding the best performance. In higher dimensions with the longest sequence setting, the benefits of the balanced forward KL loss, with its density minimization effect in areas of lower utility, are more evident. In addition, we note that exponential regularization, corresponding to assuming an exponential dot-product kernel for the RKHS feature space of the model (see 2), allowed for the best performance in higher dimensions. Lastly, in Figure 6 (appendix), we present an ablation study on the batch size setting $B$, showing monotonic improvements, especially for large $B\geq 32$.
(a) Stability
(b) SASA
Figure 3: Batch diversity scores per round on the FoldX protein optimization tasks.
For our final set of experiments we present two real protein optimization tasks. These experiments have been adapted from stanton2022accelerating where the aims are to maximize the stability and solvent accessible surface (SASA) of the proteins, respectively. The black-box is the FoldX molecular simulation software (Schymkowitz2005foldx), and is wrapped by poli (Gonzalez-Duque2024poli). We chose the mRouge red fluorescent protein ($M=228$) as the base protein for the tasks. Both tasks were given $T=20$ rounds, a batch size of $B=64$, and an initial training set of $|{\mathcal{{D}}}_{0}|=88$ as a subset from stanton2022accelerating. Results are summarized in 1(b) for stability and 1(c) for SASA. All variants of GenBO find the stability task challenging, along with the LaMBO-2 and random baselines. CbAS and especially VSD are better able to stabilize this protein. As shown by diversity scores in 3(a), which we measure by averaging the Levenshtein distance across the batch in the same way as Steinberg2025variational, algorithmic baselines with the lowest diversity yielded top performance, indicating that pure exploitation from around the starting dataset led to the highest outcomes. However, most variants of GenBO far outperform the baselines on the SASA task, and much more rapidly. We believe this task favors extrapolation away from the prior, due to the high performance of GenBO variants with uninformative prior. In contrast to the stability, the diversity scores show that increasing exploration led to better outcomes for SASA (3(b)).